OpenStudy (anonymous):

how do you put this in y=mx+b form? y=x+2/4

OpenStudy (phi):

is that \[ y=\frac{x+2}{4} \]?

OpenStudy (anonymous):

yess

OpenStudy (phi):

you can rewrite the fraction as x/4 + 2/4

OpenStudy (anonymous):

y=(1/4)x+(1/2)

OpenStudy (anonymous):

m=(1/4) b=(1/2)

OpenStudy (anonymous):

thanks so much!